OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±10) and asymptotes at y = ±5/6* x.

OpenStudy (plasmataco):

... I'll try but no garentees

OpenStudy (anonymous):

No problem! Atleast if you can try or get someone to help you too

OpenStudy (plasmataco):

It's center is at the origin so...

OpenStudy (plasmataco):

Ok... U listening?

OpenStudy (plasmataco):

What u got to do is put the x in the coordinates of the vertices in for the asymptote equation.

OpenStudy (plasmataco):

So plug 10 in to get y=50/6

OpenStudy (anonymous):

is this vertical or horiztonal though

OpenStudy (plasmataco):

.

OpenStudy (plasmataco):

Hang on.

OpenStudy (plasmataco):

Because the center is (0,0), it makes thing a lot easier

OpenStudy (anonymous):

Oh okay

OpenStudy (plasmataco):

The equation is ((X-g)/a)-((y-f)/b)=1

OpenStudy (anonymous):

okay yeah I have that

OpenStudy (plasmataco):

(G,f) is the center of the hyperbola in which case (0,0) so we can cancel those out to get (X squared)/a-(y squared)/b

OpenStudy (plasmataco):

My first equation is wrong. I forgot the power 2 after(x-g) and (y-f) sry 😓

OpenStudy (anonymous):

okay yeah I was gonna correct that

OpenStudy (plasmataco):

A is the x part of the vertice so 10. B is what we solved earlier with the asymptote.

OpenStudy (anonymous):

wait what is b?

OpenStudy (plasmataco):

So we get (x squared)/10-(y squared times 6)/50=1

OpenStudy (plasmataco):

The 6th comment

OpenStudy (plasmataco):

No wait the answer is (x/10) squared-(6y/50) squared=1

OpenStudy (plasmataco):

Forgot another pair of squares at the denominator

OpenStudy (plasmataco):

Here's the link I used

OpenStudy (plasmataco):

http://www.endmemo.com/geometry/hyperbola.php

OpenStudy (plasmataco):

@PHUNISH ?

OpenStudy (plasmataco):

Ok well. I gotta go sry

OpenStudy (plasmataco):

I hoped I helped

OpenStudy (anonymous):

Yeah sorry I been trying to figure this out with my notes but you helped so much more thank you so much